Thirty people, who are of the same age and the same health status, are

insured with the same insurance company. Using life tables, the company

estimates that the probability for a randomly chosen person among these 30

to be alive in 15 years from now is 0:8.

(i) What is the probability that not all 30 people will be alive in 15 years time?

(ii) What is the probability that at least one person will be alive?


1
Expert's answer
2022-05-09T18:32:43-0400

Let "X=" the number of person who will be alive:"X\sim Bin(n, p)."

Given "n=30, p=0.8, q=1-p=0.2."

(i)

"P(X<30)=1-P(X=30)"

"=1-\dbinom{30}{0}(0.8)^{30}(0.1)^{30-30}"

"=0.99876205996"

(ii)


"P(X\ge1)=1-P(X=0)"

"=1-\dbinom{30}{0}(0.8)^{0}(0.2)^{30-0}"

"> 0.999999"